# homogeneous_khovanskii.bib

### v1

@@ 5.2. Comparison to Yamagishi’s algorithm
By [Yam18, Corollary 4.13], a finite $\operatorname{Ab}(G)^{\vee}$/̄homogeneous MUVAK basis of $\mathbb{C}[V]^{[G,G]}$ with respect to $v_{1},\dots,v_{m}$ exists. Hence, by Proposition 3.16, Algorithm 3 is applicable to compute such a basis.
@@
Notice that $\mathbb{C}[V]^{[G,G]}$ is a standard graded ring and that the action of $\operatorname{Ab}(G)$ on $\mathbb{C}[V]^{[G,G]}$ is linear, so preserves the standard degree. In consequence, the grading by $\operatorname{Ab}(G)^{\vee}$ is well/behaved in the sense that we may decompose any $\operatorname{Ab}(G)^{\vee}$/̄homogeneous polynomial $f\in\mathbb{C}[V]^{[G,G]}$ as a sum of standard homogeneous polynomials which are $\operatorname{Ab}(G)^{\vee}$/̄homogeneous as well.
@@
In [Yam18], Yamagishi gives an algorithm that computes a $\operatorname{Ab}(G)^{\vee}$/̄homogeneous MUVAK basis of $\mathbb{C}[V]^{[G,G]}$ with respect to $v_{1},\dots,v_{m}$. In broad strokes, Yamagishi’s algorithm first adds generators until the set of generators is a MUVAK basis for every valuation $v_{i}$ individually, that is, we have generators $f_{1},\dots,f_{k}\in\mathbb{C}[V]^{[G,G]}$ such that for all $f\in\mathbb{C}[V]^{[G,G]}$ there are polynomials $h_{1},\dots,h_{m}\in\mathbb{C}[X_{1},\dots,X_{k}]$ with $h_{i}(f_{1},\dots,f_{k})=f$ and $v_{i}(t(f_{1},\dots,f_{k}))\geq v_{i}(f)$ for all $t\in\operatorname{Supp}(h_{i})$ and $1\leq i\leq m$. Afterwards, the algorithm considers pairs, triples, and so on, of valuations to ensure that eventually one has a MUVAK basis with respect to $v_{1},\dots,v_{m}$, that is, that for all $f\in\mathbb{C}[V]^{[G,G]}$ there is $h\in\mathbb{C}[X_{1},\dots,X_{k}]$ with $h(f_{1},\dots,f_{k})=f$ and $v_{i}(t(f_{1},\dots,f_{k}))\geq v_{i}(f)$ for all $t\in\operatorname{Supp}(h)$ and $1\leq i\leq m$.
We implemented the algorithms in this paper in the computer algebra system OSCAR [Osc, DEFHJ24]. We also implemented Yamagishi’s algorithm as part of [Sch23] and this is to our knowledge the only available implementation of the algorithm in [Yam18]. We now compare the performance of the different algorithms. For this, we ran the algorithms with certain symplectic reflection groups $G\leq\operatorname{GL}(V)$ [Coh80]. The corresponding linear quotients $V/G$ are of particular interest as they are examples of symplectic singularities [Bea00], see [Sch23, Chapter 2] for an overview and more references. A large class of symplectic reflection groups is related to complex reflection groups: given a complex reflection group $H\leq\operatorname{GL}(W)$, where $W$ is a finite/dimensional complex vector space, the group
Table 1. Runtimes of computing a $\operatorname{Ab}(G)^{\vee}$/̄homogeneous MUVAK basis of $\mathbb{C}[V]^{[G,G]}$ with respect to the valuations arising from McKay correspondence (all times in seconds)
@@
$G_{12}^{\circledast}$ 1 $>36000$ $415$ $>36000$
@@ Appendix A A generalization of Bayer’s method
In Algorithm 3, we need to compute the homogenization of an ideal with respect to non/negative weights. This involves the computation of a saturation, for which a naive approach requires potentially several expensive Gröbner basis computations. In the following, we describe how one may adapt an algorithm known as ‘Bayer’s method’ to compute such a saturation, and hence the homogenization, efficiently. The results of this appendix were already published as part of [Sch23]; we repeat them here for the reader’s convenience.
We require some notions from the theory of Gröbner bases (or standard bases) and refer to [GP08, Chapter 1] for the basic definitions. In particular, if $>$ is a monomial ordering on a polynomial ring $K[X_{1},\dots,X_{k}]$, then following [GP08, Definition 1.2.1] we allow that $X_{i}<1$, which is occasionally excluded in the definition of monomial orderings. We also adopt the terminology of only speaking of ‘Gröbner bases’ if the ordering is global and of ‘standard bases’ in general, see [GP08, Definition 1.6.1].
[…]
If $w_{i}\neq 0$ for all $1\leq i\leq k$, there is a quite simple method to compute $I^{\mathrm{hom}}$ that only involves the computation of a Gröbner basis of $I$ with respect to the weighted degree ordering defined by $\mathbf{w}$, see [GP08, Exercise 1.7.5]. However, although the weights in our application are non/negative, they might in general be zero, so they do not give rise to a total ordering on the set of monomials and we cannot make use of this approach.
@@
This is a matrix of full rank as $w_{k}\neq 0$ and hence induces a monomial ordering $>_{M}$ on the monomials of $K[X_{1},\dots,X_{k},t]$ by multiplying the exponent vectors by $M$ (from the left) and then using the lexicographic ordering on $\mathbb{Z}^{k+1}$, see [GP08, Remark 1.2.7].
@@
Point (d) is the direct generalization of the above mentioned ‘core idea’ for Bayer’s method: for a $\deg_{\mathbf{w}}$/̄homogeneous polynomial $f$, we have $t\mid\operatorname{LT}_{>_{M}}(f)$ if and only if $t\mid f$, where $\operatorname{LT}_{>_{M}}$ is the leading term with respect to $>_{M}$. It follows from point (a) that $>_{M}$ is global with respect to the variables $X_{1},\dots,X_{k}$ and this also extends to the variable $t$, if $\deg_{\mathbf{w}}(t)=1$, that is, if we homogenize positively, by point (b). However, if $\deg_{\mathbf{w}}(t)=-1$, then $t<_{M}1$, so the ordering $>_{M}$ is local with respect to $t$. This second case is more challenging: in order to speak about a standard basis of $\tilde{I}$ with respect to $>_{M}$, we have to consider the extension of $\tilde{I}$ to the localization
@@
In case $\deg_{\mathbf{w}}(t)=1$, the ordering $>_{M}$ is global, so we have $S=K^{\times}$ by [GP08, p. 39].
@@
With the above notation, let $J\trianglelefteq K[X_{1},\dots,X_{k},t]$ be a $\deg_{\mathbf{w}}$/̄homogeneous ideal. Then we have
@@
Let $J\trianglelefteq K[X_{1},\dots,X_{k},t]$ be a $\deg_{\mathbf{w}}$/̄homogeneous ideal and let $g_{1},\dots,g_{s}\in K[X_{1},\dots,X_{k},t]$ be a standard basis of $S^{-1}J$ with respect to the monomial ordering $>_{M}$. Write $g_{i}=t^{m_{i}}g^{\prime}_{i}$ for $g^{\prime}_{i}\in K[X_{1},\dots,X_{k},t]$ with $t\nmid g^{\prime}_{i}$. Then $g^{\prime}_{1},\dots,g^{\prime}_{s}$ generate $J:\langle t\rangle^{\infty}$.
@@
We may assume that the $g_{i}$ are $\deg_{\mathbf{w}}$/̄homogeneous, as replacing $g_{i}$ by their homogeneous parts does not change the standard basis property. Therefore, $t\nmid\operatorname{LT}_{>_{M}}(g^{\prime}_{i})$ by choice of $g^{\prime}_{i}$ and the properties of $>_{M}$. So, $\operatorname{LT}_{>_{M}}(g^{\prime}_{i})\mid\operatorname{LT}_{>_{M}}(g^{\prime})$ and this proves that $g^{\prime}_{1},\dots,g^{\prime}_{s}$ is a standard basis of $S^{-1}\big{(}J:\langle t\rangle^{\infty}\big{)}$, so in particular a generating system [GP08, Lemma 1.6.7 (3)]. But then $g^{\prime}_{1},\dots,g^{\prime}_{s}$ generate $J:\langle t\rangle^{\infty}$ by Lemma A.2 again. ∎
In conclusion, to compute $I^{\mathrm{hom}}=\tilde{I}:\langle t\rangle^{\infty}$ we need to compute a standard basis for $\tilde{I}$ with respect to $>_{M}$ and divide the elements by $t$. This only involves the computation of one standard basis and proved to be quite efficient in practice compared with the computation of the saturation via iterated quotients. To compute the homogenization with respect to several gradings $\deg_{1},\dots,\deg_{m}$ as in Algorithm 3, we may use this approach iteratively.
[…]

### v2

@@ 5.2. Comparison to Yamagishi’s algorithm
By [Yam18, Corollary 4.13], a finite $\operatorname{Ab}(G)^{\vee}$-homogeneous MUVAK basis of $\mathbb{C}[V]^{[G,G]}$ with respect to $v_{1},\dots,v_{m}$ exists. Hence, by Proposition 3.21, Algorithm 3 is applicable to compute such a basis.
@@
Notice that $\mathbb{C}[V]^{[G,G]}$ is a standard graded ring and that the action of $\operatorname{Ab}(G)$ on $\mathbb{C}[V]^{[G,G]}$ is linear, so preserves the standard degree. In consequence, the grading by $\operatorname{Ab}(G)^{\vee}$ is well-behaved in the sense that we may decompose any $\operatorname{Ab}(G)^{\vee}$-homogeneous polynomial $f\in\mathbb{C}[V]^{[G,G]}$ as a sum of standard homogeneous polynomials which are $\operatorname{Ab}(G)^{\vee}$-homogeneous as well.
@@
In [Yam18], Yamagishi gives an algorithm that computes a $\operatorname{Ab}(G)^{\vee}$-homogeneous MUVAK basis of $\mathbb{C}[V]^{[G,G]}$ with respect to $v_{1},\dots,v_{m}$. In broad strokes, Yamagishi’s algorithm first adds generators until the set of generators is a MUVAK basis for every valuation $v_{i}$ individually, that is, we have generators $f_{1},\dots,f_{k}\in\mathbb{C}[V]^{[G,G]}$ such that for all $f\in\mathbb{C}[V]^{[G,G]}$ there are polynomials $h_{1},\dots,h_{m}\in\mathbb{C}[X_{1},\dots,X_{k}]$ with $h_{i}(f_{1},\dots,f_{k})=f$ and $v_{i}(t(f_{1},\dots,f_{k}))\geq v_{i}(f)$ for all $t\in\operatorname{Supp}(h_{i})$ and $1\leq i\leq m$. Afterwards, the algorithm considers pairs, triples, and so on, of valuations to ensure that eventually one has a MUVAK basis with respect to $v_{1},\dots,v_{m}$, that is, that for all $f\in\mathbb{C}[V]^{[G,G]}$ there is $h\in\mathbb{C}[X_{1},\dots,X_{k}]$ with $h(f_{1},\dots,f_{k})=f$ and $v_{i}(t(f_{1},\dots,f_{k}))\geq v_{i}(f)$ for all $t\in\operatorname{Supp}(h)$ and $1\leq i\leq m$.
Table 1. Runtimes of computing a $\operatorname{Ab}(G)^{\vee}$-homogeneous MUVAK basis of $\mathbb{C}[V]^{[G,G]}$ with respect to the valuations arising from McKay correspondence (all times in seconds)
@@
$G_{12}^{\circledast}$ 1 >36000>36000 $415$ >36000>36000
We implemented the algorithms in this paper in the computer algebra system OSCAR [Osc25, Dec+25]. This implementation is freely available online at
https://gitlab.com/math5724907/homogeneouskhovanskii.
We also implemented Yamagishi’s algorithm as part of [Sch23] and this is to our knowledge the only available implementation of the algorithm in [Yam18]. We now compare the performance of the different algorithms. For this, we ran the algorithms with certain symplectic reflection groups $G\leq\operatorname{GL}(V)$ [Coh80]. The corresponding linear quotients $V/G$ are of particular interest as they are examples of symplectic singularities [Bea00], see [Sch23, Chapter 2] for an overview and more references. A large class of symplectic reflection groups is related to complex reflection groups: given a complex reflection group $H\leq\operatorname{GL}(W)$, where $W$ is a finite-dimensional complex vector space, the group
@@ Appendix A A generalization of Bayer’s method
In Algorithm 3, we need to compute the homogenization of an ideal with respect to non-negative weights. This involves the computation of a saturation, for which a naive approach requires potentially several expensive Gröbner basis computations. In the following, we describe how one may adapt an algorithm known as ‘Bayer’s method’ to compute such a saturation, and hence the homogenization, efficiently. The results of this appendix were already published as part of [Sch23]; we repeat them here for the reader’s convenience.
We require some notions from the theory of Gröbner bases (or standard bases) and refer to [GP08, Chapter 1] for the basic definitions. In particular, if >> is a monomial ordering on a polynomial ring $K[X_{1},\dots,X_{k}]$, then following [GP08, Definition 1.2.1] we allow that $X_{i}<1$, which is occasionally excluded in the definition of monomial orderings. We also adopt the terminology of only speaking of ‘Gröbner bases’ if the ordering is global and of ‘standard bases’ in general, see [GP08, Definition 1.6.1].
For the following discussion, let $K[X_{1},\dots,X_{k}]$ be graded by an integral weight vector $\mathbf{w}=(w_{1},\dots,w_{k})\in\mathbb{Z}^{k}_{\geq 0}$ via $\deg_{\mathbf{w}}(X_{i})=w_{i}$ and let $I\trianglelefteq K[X_{1},\dots,X_{k}]$ be an ideal. We add an additional variable $t$ to $K[X_{1},\dots,X_{k}]$ and want to compute the homogenization $I^{\mathrm{hom}}$ of $I$ with respect to the grading $\deg_{\mathbf{w}}$ and the variable $t$. Depending on whether we want to homogenize ‘positively’ or ‘negatively’, we set $\deg_{\mathbf{w}}(t)\coloneqq 1$ or $\deg_{\mathbf{w}}(t)\coloneqq-1$, respectively.
If $w_{i}\neq 0$ for all $1\leq i\leq k$, there is a quite simple method to compute $I^{\mathrm{hom}}$ that only involves the computation of a Gröbner basis of $I$ with respect to the weighted degree ordering defined by $\mathbf{w}$, see [GP08, Exercise 1.7.5]. However, although the weights in our application are non-negative, they might in general be zero, so they do not give rise to a total ordering on the set of monomials and we cannot make use of this approach.
@@
This is a matrix of full rank as $w_{k}\neq 0$ and hence induces a monomial ordering $_{M}$ on the monomials of $K[X_{1},\dots,X_{k},t]$ by multiplying the exponent vectors by $M$ (from the left) and then using the lexicographic ordering on $\mathbb{Z}^{k+1}$, see [GP08, Remark 1.2.7].
@@
Point (d) is the direct generalization of the above mentioned ‘core idea’ for Bayer’s method: for a $\deg_{\mathbf{w}}$-homogeneous polynomial $f$, we have $t\mid\operatorname{LT}_{>_{M}}(f)$ if and only if $t\mid f$, where $\operatorname{LT}_{>_{M}}$ is the leading term with respect to $_{M}$. It follows from point (a) that $_{M}$ is global with respect to the variables $X_{1},\dots,X_{k}$ and this also extends to the variable $t$, if $\deg_{\mathbf{w}}(t)=1$, that is, if we homogenize positively, by point (b). However, if $\deg_{\mathbf{w}}(t)=-1$, then $t<_{M}1$, so the ordering $_{M}$ is local with respect to $t$. This second case is more challenging: in order to speak about a standard basis of $\tilde{I}$ with respect to $_{M}$, we have to consider the extension of $\tilde{I}$ to the localization
@@
In case $\deg_{\mathbf{w}}(t)=1$, the ordering $_{M}$ is global, so we have $S=K^{\times}$ by [GP08, p. 39].
@@
With the above notation, let $J\trianglelefteq K[X_{1},\dots,X_{k},t]$ be a $\deg_{\mathbf{w}}$-homogeneous ideal. Then we have
@@
Let $J\trianglelefteq K[X_{1},\dots,X_{k},t]$ be a $\deg_{\mathbf{w}}$-homogeneous ideal and let $g_{1},\dots,g_{s}\in K[X_{1},\dots,X_{k},t]$ be a standard basis of $S^{-1}J$ with respect to the monomial ordering $_{M}$. Write $g_{i}=t^{m_{i}}g^{\prime}_{i}$ for $g^{\prime}_{i}\in K[X_{1},\dots,X_{k},t]$ with $t\nmid g^{\prime}_{i}$. Then $g^{\prime}_{1},\dots,g^{\prime}_{s}$ generate $J:\langle t\rangle^{\infty}$.
@@
We may assume that the $g_{i}$ are $\deg_{\mathbf{w}}$-homogeneous, as replacing $g_{i}$ by their homogeneous parts does not change the standard basis property. Therefore, $t\nmid\operatorname{LT}_{>_{M}}(g^{\prime}_{i})$ by choice of $g^{\prime}_{i}$ and the properties of $_{M}$. So, $\operatorname{LT}_{>_{M}}(g^{\prime}_{i})\mid\operatorname{LT}_{>_{M}}(g^{\prime})$ and this proves that $g^{\prime}_{1},\dots,g^{\prime}_{s}$ is a standard basis of $S^{-1}\big(J:\langle t\rangle^{\infty}\big)$, so in particular a generating system [GP08, Lemma 1.6.7 (3)]. But then $g^{\prime}_{1},\dots,g^{\prime}_{s}$ generate $J:\langle t\rangle^{\infty}$ by Lemma A.2 again. ∎
In conclusion, to compute $I^{\mathrm{hom}}=\tilde{I}:\langle t\rangle^{\infty}$ we need to compute a standard basis for $\tilde{I}$ with respect to $_{M}$ and divide the elements by $t$. This only involves the computation of one standard basis and proved to be quite efficient in practice compared with the computation of the saturation via iterated quotients. To compute the homogenization with respect to several gradings $\deg_{1},\dots,\deg_{m}$ as in Algorithm 3, we may use this approach iteratively.